OpenStudy (kkutie7):

A spring has a natural length of 30 cm. If a 24-N force is required to keep it stretched to a length of 42 cm, how much work W is required to stretch it from 30 cm to 36 cm? (Round your answer to two decimal places.)

OpenStudy (kkutie7):

Work=force*distance

OpenStudy (kkutie7):

i need to find an equation for right?

OpenStudy (kkutie7):

*work

OpenStudy (zephyr141):

is that 24 N or 24 minus N?

OpenStudy (kkutie7):

24 newtons

OpenStudy (zephyr141):

ok. i just see a minus there. had to clarify

OpenStudy (kkutie7):

so any ideas

OpenStudy (kkutie7):

I'm actually supposed to find an equation for work and put it into an integral from 30 to 36 cm i just dont know how to start

OpenStudy (kkutie7):

@kirbykirby

OpenStudy (aum):

F = k * x 24 N force to stretch it by (42-30 = 12) cm. k = F / x = 24/12 = 2 N / cm W = Force * Distance dW = kx * dx \[ W = \int_{x_1}^{x_2}kxdx = \int_{30}^{36}2xdx = ? \]

OpenStudy (kkutie7):

396?

OpenStudy (aum):

Yes, but you may have to convert units so that he answer is is Newton-meters instead of 396 Newton-centimeters.

OpenStudy (kkutie7):

it needs to be in Jouls

OpenStudy (aum):

Yes, 396 Newton-Centimeters or 3.96 Newton-Meters = 3.96 Joules.

OpenStudy (kkutie7):

ok great